LUCI Technical Observing, prep for Partner Science:
UT 2011 MMM DD

Observer:
Telescope Operator:
Instrument Support:
Telescope Support:
SW support:

Summary

Preparations

Details

Restart Task List, see: http://wiki.lbto.org/twiki/bin/view/Instrumentation/RecalibrationOfLUCIAfterInstall

Observing Plan (copy and paste this to the nightly logs)

LUCI1 Daily functional checkout:
  • Run /home/engineer/bin/showSupervisorStatus.sh, check if OK
  • Check LUCI1 Web-IO Control, verify that ROE and MCE are powered
  • Check is calibration unit is in the out position
  • Open UI and check that telescope data are being updated
    • If not, restart Telescope Service through SW Control Center
  • Verify data are moving to archive

LUCI1 Weekly Operations: (do prior to each partner science block)
  • Check darks (see: LBTtools/LUCI/readnoise script)
    • O2DCR 2.0s (usually around 11.6 e-, 12.6 for detector #2)
    • MER10 10.0s (usually around 5.5 e-, 7.75 for detector #2)
  • Check/adjust Pupil and Field Stop Alignment
  • Check /data/luci, make room if necessary
    • login as luci@lucifer (must be the owner to delete files)
    • Before deleting data from lucifer, check /Repository/
    • If discrepancies are found:
      • Check /newdata/indir, /newdata/warning
      • Check the archive web interface
    • If necessary: cp /archive

LUCI1 Monthly Operations:
  • Check system time vs our GPS, at cmd line execute: "ntpdate -q gps0.mountain.lbto.org"
  • Check mask tables after exchange on the Mask Config GUI
    • Make sure the column 'in use' does not have a check mark where there is no mask.
    • Make sure there are no blank spaces before the mask name
    • LMS files:
      • Make sure the path name is correct
      • Verify the lms files exist in this directory on the lucifer computer.
  • Shutdown and restart LUCIFER Control SW after mask exchange
  • Take sieve mask images in N3.75 and N1.8 to check focus gradients
  • Move calibration unit in
    • Check that lamps work
    • Check gratings/tilts
    • Take an imaging flat and check levels, dirt on window
    • Move calibration unit out

Other Tests:

Closed dome SW tests, if requested
  • Test sound playing through IIF (write iraf task to do this)

Twilight flats in NB, BB filters (cycle through all over time)
  • Need to set up a records page on the wiki to keep track

Correct initial collimation/pointing

Open dome SW tests, if requested
  • Optical pointing/co-pointing/range balancing algorithm, also using IRTC

Take set of linearization test data (photometric weather only!)
  • Preset to any Persson standard using the catalog interface
  • source the shell script ~/LBTO/SciencePrep/LinTestData/lintest
  • There is a README file in this directory describing the test
  • Should take ~25 minutes to run.

Test any new template observing scripts prior to release

Run science-type dice5 pattern to check that things still work
  • Preset to available BS91XX star from UI/Catalog
  • Execute script: ~/LBTO/SciencePrep/dice5_10x (~1 hr)
  • Any other source with a known ABS dither pattern is OK
  • Calculate rms/position, scale & rotation of pattern

Repeat check of dithering along slit (20x 2' dither)
  • Preset to available BS91XX star from UI/Catalog
  • Execute script: ~/LBTO/SciencePrep/slitdither.acq
  • Align BS source to center of 2" longslit
  • Execute script: ~/LBTO/SciencePrep/slitdither.sci (~30 min)
  • Calculate AGw alignment to slit

Collect open-loop tracking data to check stability
  • Preset ACTIVE on BS91XX to collimate
  • Preset track to same source
  • Collect long series of 30x10s exposures, saved NORMAL

Repeat measurement of focal plane scale, orientation
  • Any low-latitude field with 20+ stars in USNO is OK
  • Random fields OK, or some BS91xx fields with lots of stars
  • Execute script: ~/LBTO/SciencePrep/imagepair

Additional work on M1/M2 range-balancing collimation

Field aberration data collection

Check existing collimation model
  • Collimate at mid-EL
  • Check at low EL w/TRACK preset
  • Check at high EL w/TRACK preset

Check existing pointing model
  • Implement new collimation model first, if necessary
  • Correct pointing at mid-EL
  • Check a few points
  • Stable conditions!
  • It would be useful to take a set of 12 points if possible
  • Use the "pointing" mask for focal plane measurements.

Photometric zero point and extinction monitoring

Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
  • Check to make sure offsets still happen in a reasonable manner

Improve map of AGw patrol field limits at high field angle
  • Requires set up work on a stone field

Gather NB illumination correction data (BrG done on 20100606)
  • Needs photometric weather
  • See ~/LBTO/SciencePrep/Illumination/README.txt for info

Field Check - verification of WFS algorithm

Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)

Test dithering between extreme slits on MOS mask + telluric
  • Manual needs 0.59836mm/" scale for the dXdY calculation
  • The how-to set up telluric observations needs updating for clarity.

Investigate variable imaging FF gradients
  • Take data while the telescope is offset with respect to the enclosure
  • Take data at PA=X, PA=X+180

Tests that have been satisfactorily completed and do not need repetition:

Re-test updatePointingReference functionality (high priority) (June 2011)
  • See description wiki page here
  • Scripts are in ~/LBTO/SciencePrep/zeroBase
  • Manually preset to BS91XX star at non-cardinal PA
  • Execute appropriate offset script(s)
  • Ask OSA to "absorb" in the appropriate coord sys.
  • Execute appropriate dice5 script(s)
  • 9 blocks x ~5 minutes each

Test WCS/ROTCEN Coordinates issue
  • Preset to any Persson standard, allow to collimate
  • cd ~/LBTO/SciencePrep/WCStest (there is a README file)
  • ./wcstest.sh (executes setup then the four dice5 scripts)
  • Should take less than 30 minutes.
  • Result: CRVAL1/2 in header walks off where it should be...next step in PCS

Test defocus on AGw by 13mm, Johanna's acq script
  • Need decent (<1") seeing.
  • Ask OSA to do the following on obs1:
    • Edit /lbt/tcs/current/etc/GCS/AIP_L.cfg
    • change "focus" to be 44 instead of 31, save
    • selectAGW left AIP_L.cfg
  • Preset to target field:
    • ~/LBTO/2011.02/WASP12b_acquisition_longslit (06:30:32.7907 +29:40:20.440)
    • Use any Persson standard if WASP12b is not up
  • Allow AGw to re-collimate (wait for collimation ON)
  • Can execute the corresponding science script if on WASP12b (at least 1 pair of spectra)
  • Ask OSA to do the following on obs1:
    • Edit /lbt/tcs/current/etc/GCS/AIP_L.cfg
    • change "focus" back to 31, save
    • selectAGW left AIP_L.cfg

DThompson notes (not part of plan above)

  • Investigate illumination changes in twilight flats, build library of flats
  • Improve collimation
  • Improve pointing
  • Work on co-pointing SX and DX
  • Measure/correct field aberration issues (aligns optical axes of M1, M2)
  • Monitor orientation of detector on sky
  • Diagnose a possible WCS issue (reported last week by INAF)
  • Monitor/correct AGw transformation
  • Investigate photometric gradients (BB=5%, NB=25%)
  • Test new functionality (TCS and/or LUCI)
  • Test various approaches to scripting
  • Monitor photometric zero points

# * Take a set of (monocular) range-balanced collimation data # * Stable conditions! # * implement this collimation model

# * Take a set of pointing data using range-balanced collimation model # * Stable conditions!

# * Collect field aberration data

# * Collect through-focus sequence for Andrew - reasonably dense star field. # * Try Z4 of -1000, -500. 0, 500, 1000 nm # * Also check the AGw-LUCIFER focus in cold weather for temp term?

# * Collect astrometric field data to check AGw calibration for monitoring

-- DavidThompson - 07 Apr 2010
Topic revision: r55 - 07 Oct 2013, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback